OpenStudy (anonymous):

Solve for the variable: -12x + 12 = -24

jimthompson5910 (jim_thompson5910):

First step?

OpenStudy (anonymous):

-12x = -24 - 12 Multiply by (-1) both sides, 12x = 24 + 12 or, x = 3 is the answer...